# How to Bridge the Sim-to-Real Gap in Digital Twin-Aided Telecommunication Networks

### v2

@@ I Introduction
How can we bridge the sim-to-real gap to ensure that AI models trained with DTs remain reliable in real deployments? As illustrated in Fig. 2, this article explores two complementary strategies:
@@
As a final remark, we emphasize that this paper focuses on leveraging DTs for data augmentation in AI model training. While this represents an important use case, as highlighted by O-RAN [oran2024dtran, li2025generative], DTs can also enable real-time operational tasks, including, for instance, channel and link quality prediction [DTrev]. Such real-time applications require continuous synchronization mechanisms to manage the real-to-sim gap at fine temporal granularities [hashash2022edge]. In contrast, our work addresses DT calibration at a non-real-time scale, focusing on training AI models using DTs that capture long-term environmental statistics and operating conditions.
@@
Figure 2: As detailed in this paper, the sim-to-real gap can be partially bridged via three main complementary strategies: (a) Digital twin calibration (Sec. II), which uses real-world measurements to directly improve the adherence of the DT to the real world; (b)-(c) Robust AI training, which accounts for the residual sim-to-real gap by modeling the uncertainty on the true environment via Bayesian inference (panel (b), Sec. III), or by correcting the bias caused by the sim-to-real gap on the training objective (panel (c), Sec. IV). (Dashed lines indicate DT-generated synthetic data, while solid lines represent real-world measurements.)
@@ II-B Phase Error-Aware Calibration
Calibrating material parameters is useful, but it cannot fully resolve biases caused by a misspecified geometry of the objects in the scene. For instance, as seen in Fig. 2(a), minor errors in building shapes or additional objects not present in the DT can cause significant phase errors in the simulated propagation paths. In fact, even a small displacement of an object—of the order of a fraction of the wavelength—can cause the phase of a reflected path to shift, leading to constructive or destructive interference patterns that do not align with reality [ruah2024calibrating]. Owing to these unaccounted phase errors, traditional calibration methods that only tweak material values $\rho$ may fail to produce an accurate simulator.
To tackle this issue, a promising approach is to devise calibration schemes that are aware of the presence of phase errors caused by a sim-to-real gap in the geometry of the environment. The core idea behind phase error-aware calibration, as proposed in [ruah2024calibrating], is to iteratively estimate and correct phase errors in the simulated channel. In this method, the DT’s prediction is augmented with per-path phase error terms that account for unmodeled geometry nuances. An expectation-maximization (EM) algorithm is then used: in the E-step, the algorithm infers the likely phase errors for each propagation path by comparing the simulation with real measurements; while the M-step updates the simulator parameters $\rho$ to better fit the observed data, effectively absorbing the phase errors into the model.
@@
By disregarding phase errors, phase-oblivious calibration is seen to be unable to provide satisfactory results across all bandwidth levels. Uniform phase error calibration yields improved results as the bandwidth levels approach the delay-resolution needed to separate the contributions of the interfering paths. Finally, phase error-aware calibration can dramatically reduce the power prediction error as compared to the two baseline methods, even at lower bandwidth levels. This demonstrates that explicitly accounting for local phase error during calibration can significantly improve the reliability of the ray tracer’s estimates. This advantage is especially pronounced in scenarios with a rich multipath structure, in which small geometry differences would otherwise lead to large interference errors.
@@ III Modeling the Sim-to-Real Gap on the Environment
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [simeone2022machine, Ch. 12]. It should be emphasized that, while these approaches improve robustness and reliability, they do come at the cost of increased computational complexity: sampling multiple environments, training across a population of models, and maintaining distributions over parameters may introduce significant overhead. Nonetheless, in many critical applications where reliability and safety are paramount, this additional cost often offers a worthwhile trade-off.
@@ IV Modeling the Sim-to-Real Gap on the Training Objective
Another approach to address the residual sim-to-real gap is to leave the DT as is, but modify the training algorithm to compensate for the DT’s bias. In this regard, a recently developed framework called PPI [angelopoulos2023prediction] offers a powerful tool for training AI models when the amount of real world data is too small to reach the desired level of accuracy. PPI originates from the idea of doubly robust estimation in statistics, and it supports the use of abundant biased data from a simulator, together with a small amount of unbiased data from the real world, to produce unbiased learning outcomes.
@@ IV-A Semi-Supervised Learning via Prediction-Powered Inference
In contrast, as illustrated in Fig. 2(c), PPI uses the real data not to train the model directly, but to estimate the bias of a training loss estimate based on synthetic data. It then adjusts the model training objective to remove this bias. Essentially, the real data provides a yardstick for how far off the synthetic training loss is from the real training loss. PPI is appealing for telecommunication AI problems because obtaining even a small validation set of real data is often feasible through field tests or pilot deployments, whereas gathering a comprehensive training set is not [sifaou2024semi].
@@ V Conclusion
As AI continues to scale, its promise for revolutionizing network automation and optimization is tempered by the realities of deployment in complex, data-scarce telecommunication environments. DTs have emerged as a vital tool to generate the data needed to train and test AI solutions virtually. By calibrating DTs with real-world data, and by explicitly modeling residual sim-to-real gaps through Bayesian and PPI techniques, engineers can enhance the performance and reliability of DT-aided telecommunication networks.
@@
Clement Ruah (Member, IEEE) is a Research Associate at King’s Laboratory for Intelligent Computing. He received his Ph.D. degree in Information Engineering from King’s College London in 2025. His research focuses on reliable machine learning and mixing simulation-based approaches with uncertainty-aware data-driven solutions.
Houssem Sifaou (Member, IEEE) received the M.S. and Ph.D. degrees in Electrical Engineering from the King Abdullah University of Science and Technology (KAUST) in 2016 and 2021, respectively. From 2021 to 2023, he served as a Research Associate at Imperial College London. He is currently a Research Associate at King’s College London. His research lies at the intersection of machine learning and wireless communications.
Osvaldo Simeone (Fellow, IEEE) is a Professor of Information Engineering at King’s College London, where he co-directs the Centre for Intelligent Information Processing Systems and leads the King’s Communications, Learning and Information Processing lab. He is also a Visiting Professor at Aalborg University. He has received multiple best paper and tutorial awards, and his research has been supported by the NSF, ERC, EPSRC, European Commission, ESA, and industry. Prof. Simeone is a Fellow of the IEEE, IET, and EPSRC, and the author of two textbooks published by Cambridge University Press.
Bashir M. Al-Hashimi worked in the electronics design industry for eight years prior to embarking on an academic career with the School of Electronics and Computer Science, University of Southampton, in 1999. He is currently Vice President (Research & Innovation) at King’s College London and co-Director of the King’s Institute for Artificial Intelligence. He has successfully led a number of large-scale interdisciplinary research programmers funded by the EPSRC and industry, supervised 40 Ph.D. students to successful completion, published 380 refereed technical papers, and authored or coauthored seven books.

### v3

@@ I Introduction
How can we bridge the sim-to-real gap to ensure that AI models trained with DTs remain reliable in real deployments? As illustrated in Fig. 2, this article explores three complementary strategies:
@@
The three solutions described above are orthogonal and complementary. Accordingly, they can be applied on the same system in order to reap their distinct benefits. For instance, calibrating the DT reduces the need for mitigating the sim-to-real gap; while accurately modeling the sim-to-real gap on the model makes it less critical to address the sim-to-real gap on the training objective. Given that all methods are based on different principles, they are described and evaluated separately in this article. The benefits, data requirements, computational costs, and reusability across network functions (NFs) of the methods are summarized in Table I.
As a final remark, we emphasize that this paper focuses on leveraging DTs for data augmentation in AI model training. While this represents an important use case, as highlighted by O-RAN [oran2024dtran], DTs can also enable real-time operational tasks, including, for instance, channel and link quality prediction [DTrev]. Such real-time applications require continuous synchronization mechanisms to manage the real-to-sim gap at fine temporal granularities [hashash2022edge]. In contrast, our work addresses DT calibration at a non-real-time scale, focusing on training AI models using DTs that capture long-term environmental statistics and operating conditions.
@@
Figure 2: As detailed in this paper, the sim-to-real gap can be partially bridged via three main complementary strategies: (a) Digital twin calibration (Sec. II), which uses real-world measurements to directly improve the adherence of the DT to the real world; (b)-(c) Robust AI training, which accounts for the residual sim-to-real gap by modeling the uncertainty on the true environment via Bayesian inference (panel (b), Sec. III), or by correcting the bias caused by the sim-to-real gap on the training objective (panel (c), Sec. IV). (Dashed lines indicate DT-generated synthetic data, solid lines represent real-world measurements.)
TABLE I: Comparison of sim-to-real gap reduction methods
Method Benefits Data requirement Compute cost Deployment type
Calibration Improved DT Large dataset from specific deployment Training Deployment-specific, reusable across NFs
Modeling environment gap Uncertainty quantification Limited data from multiple deployments Training + sampling Deployment-specific, reusable across NFs
Modeling train ob-jective gap Debiased training objective Limited data for specific NF Training NF-specific
@@ II-B Phase Error-Aware Calibration
Calibrating material parameters is useful, but it cannot fully resolve biases caused by a misspecified geometry of the objects in the scene. For instance, as seen in the blue path of Fig. 2(a), minor errors in building shapes or additional objects not present in the DT can cause significant phase errors in the simulated propagation paths. In fact, even a small displacement of an object—of the order of a fraction of the wavelength—can cause the phase of a reflected path to shift, leading to constructive or destructive interference patterns that do not align with reality [ruah2024calibrating]. Owing to these unaccounted phase errors, traditional calibration methods that only tweak material values $\rho$ may fail to produce an accurate simulator.
To tackle this issue, a promising approach is to devise calibration schemes that are aware of the presence of phase errors caused by a sim-to-real gap in the geometry of the environment. The core idea behind phase error-aware calibration, as proposed in [ruah2024calibrating], is to iteratively estimate and correct phase errors in the simulated channel. In this method, the DT’s prediction is augmented with per-path phase error terms that account for unmodeled geometry nuances. The errors are illustrated by the phase difference of the blue path in Fig. 2(a) between the “real world” (solid line) measurement and the “uncalibrated DT” (dashed line) prediction.
The phase error estimates are used to correct the phase of the predicted paths to match real world conditions, as shown in the “calibrated DT” phase panel of Fig. 2(a). The resulting calibration algorithm alternates between phase error correction and material calibration within an expectation-maximization (EM) framework: in the E-step, the algorithm infers the likely phase errors for each propagation path by comparing the simulation with real measurements; while the M-step updates the simulator parameters $\rho$ to better fit the observed data, effectively absorbing the phase errors into the model.
@@
By disregarding phase errors, phase-oblivious calibration is seen to be unable to provide satisfactory results across all bandwidth levels. Uniform phase error calibration yields improved results as the bandwidth levels approach the delay-resolution needed to separate the contributions of the interfering paths, allowing for material parameters calibration by matching the energy of each path individually. Finally, phase error-aware calibration can dramatically reduce the power prediction error as compared to the two baseline methods, even at lower bandwidth levels. This demonstrates that explicitly accounting for local phase error during calibration can significantly improve the reliability of the ray tracer’s estimates. This advantage is especially pronounced in scenarios with a rich multipath structure, in which small geometry differences would otherwise lead to large interference errors.
@@ III Modeling the Sim-to-Real Gap on the Environment
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [11, Ch. 12].
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [simeone2022machine, Ch. 12]. The actual benefits of Bayesian learning hinge on reasonable choices for prior distribution and for the environment model, although there is a vast literature on methods that can ensure robustness to model misspecification [zecchin2023robust]. Another limitation of Bayesian learning is the increased computational complexity: sampling multiple environments, training across a population of models, and maintaining distributions over parameters may introduce significant overhead. Nonetheless, in many critical applications where reliability and safety are paramount, this additional cost often offers a worthwhile trade-off.
@@ IV Modeling the Sim-to-Real Gap on the Training Objective
Bayesian learning can help mitigate the sim-to-real gap by accounting for uncertainty at the level of the environment model. However, due to model misspecification and limited computational complexity, a residual sim-to-real gap is unavoidable in practice. This gap affects the quality of the synthetic data generated by the DT, which in turn degrades the accuracy of the training objective as an estimate of the true population loss. To address this problem, this section explores the use of PPI [angelopoulos2023prediction], which offers a powerful statistical tool for debiasing the training objective, irrespective of any level of residual sim-to-real gap on the DT.
@@ IV-A Semi-Supervised Learning via Prediction-Powered Inference
In contrast, PPI uses the real data not to train the model directly, but to estimate the bias of a training loss estimate based on synthetic data, as illustrated by the images within the parenthesis in Fig. 2(c). It then adjusts the model training objective (leftmost image in Fig. 2(c)) to remove this bias. Essentially, the real data provides a yardstick for how far off the synthetic training loss is from the real training loss. PPI is appealing for telecommunication AI problems because obtaining even a small validation set of real data is often feasible through field tests or pilot deployments, whereas gathering a comprehensive training set is not [sifaou2024semi].
@@ V Conclusion
As AI continues to scale, its promise for revolutionizing network automation and optimization is tempered by the realities of deployment in complex, data-scarce telecommunication environments. DTs have emerged as a vital tool to generate the data needed to train and test AI solutions virtually. By calibrating DTs with real-world data, and by explicitly modeling residual sim-to-real gaps through Bayesian and PPI techniques, engineers can enhance the performance and reliability of DT-aided telecommunication networks. Practitioners may decide on different combinations of the techniques presented in this work depending on the type and quality of the available data and simulator (see Table I for a summary).